if log base 2 (3) equals x-1, express log base 2 (12) in terms of x
Answer by stanbon(75887)

if log base 2 (3) equals x-1,
express log base 2 (12) in terms of x
----
12 = 2^2*3
-----
log2(12) = log2(2^2*3)
= log2(2^2) + log2(3)
= 2 + x-1
= x+1
==================
Cheers,
Stan H.